florenciam99:

Determine if the proportion 2004-04-01-01-00_files/i0260001.jpg is true. If it is not true, correct the error. A. true B. false, it should be 2004-04-01-01-00_files/i0260002.jpg C. false, it should be 2004-04-01-01-00_files/i0260003.jpg D. false, it should be 2004-04-01-01-00_files/i0260004.jpg Please select the best answer from the choices provided. A B C D Mark this and return Show Me

justasoftie:

Welcome To Qc! I wish I could tag someone to help you, but I'm unsure of anyone who is good at math other than Hero I do not believe he is on

justasoftie:

And your images are not working either

Vocaloid:

I was able to extract this from the image file the proportion is e/d = d/f answer choice B is e/d = d/h answer choice C is e/d = d/g answer choice D is f/d = d/f however, to determine which one is true we would need to know any additional context about what these variables are meant to represent